OpenStudy (anonymous):

(2b-3)^2 find the product .

OpenStudy (solomonzelman):

\[(a-b)^2=(a-b)(a-b)=a^2-2ab+b^2\]

OpenStudy (solomonzelman):

use 2b as your "a" and 3 as your "b"

OpenStudy (anonymous):

\[4b ^{2}-12b+9\]